OpenStudy (anonymous):

can someone explain x = -b/2a = -60/(2(-0.3)) = 100

jhonyy9 (jhonyy9):

x= -b/2a =10/0,1 =100

OpenStudy (anonymous):

\[\frac{-60}{2\times -.3}=\frac{-60}{-.6}=\frac{600}{6}=100\]

jhonyy9 (jhonyy9):

-60/(-0,6) =100

OpenStudy (anonymous):

ok got it thanks, it helps with the actual problem im doing